# RetrievalAttention: Accelerating Long-Context LLM Inference via Vector Retrieval

### v1

@@ Abstract
Transformer-based large Language Models (LLMs) become increasingly important in various domains. However, the quadratic time complexity of attention operation poses a significant challenge for scaling to longer contexts due to the extremely high inference latency and GPU memory consumption for caching key-value (KV) vectors. This paper proposes RetrievalAttention, a training-free approach to accelerate attention computation. To leverage the dynamic sparse property of attention, RetrievalAttention builds approximate nearest neighbor search (ANNS) indexes upon KV vectors in CPU memory and retrieves the most relevant ones via vector search during generation. Due to the out-of-distribution (OOD) between query vectors and key vectors, off-the-shelf ANNS indexes still need to scan $O(N)$ (usually 30% of all keys) data for accurate retrieval, which fails to exploit the high sparsity. RetrievalAttention first identifies the OOD challenge of ANNS-based attention, and addresses it via an attention-aware vector search algorithm that can adapt to queries and only access 1–3% of data, thus achieving a sub-linear time complexity. RetrievalAttention greatly reduces the inference cost of long-context LLM with much lower GPU memory requirements while maintaining the model accuracy. Especially, RetrievalAttention only needs 16GB GPU memory for serving 128K tokens in LLMs with 8B parameters, which is capable of generating one token in 0.188 seconds on a single NVIDIA RTX4090 (24GB).
@@ 1 Introduction
Recent transformer-based Large Language Models [1] have shown remarkable capabilities in processing long contexts. For instance, Gemini 1.5 Pro [2] has supported the context window of 10M (million) tokens. While this is promising for analyzing extensive data, longer context windows pose efficiency challenges due to the quadratic complexity of attention computation. To enhance efficiency, KV Caching, a technique that retains past Key and Value tokens, has been widely adopted to prevent redundant computations. However, KV caching-based systems face two primary issues: (a) substantial GPU memory requirements, particularly for long contexts, as exemplified by the Llama-2-7B model, which requires approximately 500GB per million tokens in FP16 format; and (b) inference latency linearly proportional to the context size, primarily due to the time needed to access cached tokens — a common issue across various computing devices, including GPUs. Therefore, reducing token access and storage costs is vital for enhancing inference efficiency.
Figure 1: Compared with existing methods, RetrievalAttention achieves the similar model accuracy as full attention but exhibits extremely low decoding latency.
@@
The solution lies in leveraging the dynamic sparsity inherent in the attention mechanism [3]. This refers to the phenomenon where each query vector significantly interacts with only a limited subset of key and value vectors, with the selection of these tokens varying dynamically based on individual queries. Numerous methods [4, 5, 6, 7, 8, 9] have been proposed to capitalize on this observation and enhance attention efficiency. However, most of these methods identify important tokens either statically [10, 11] or heuristically [6, 7, 5], leading to imprecise approximations that often result in a significant performance drop. Our analysis reveals that the Approximate Nearest Neighbor Search (ANNS) index is particularly effective in this context. When using the inner product as the distance measurement, ANNS precisely aligns with the original attention mechanism, represented as $\mathbf{q}\mathbf{K}^{T}$. It can directly identify the most critical key vectors, yielding a higher accuracy compared to previous static or heuristic methods (as illustrated in Figure 1). In this paper, we demonstrate that our ANNS-based attention (RetrievalAttention) surpasses other methods in maintaining the accuracy of full attention, particularly in long context tasks such as $\infty$-Bench [12] and RULER [13].
@@
However, integrating ANNS into attention mechanisms presents a unique challenge: the out-of-distribution (OOD) problem between query vectors and key vectors. Most ANNS engines operate under the assumption that both query and key vectors are drawn from the same distribution. To the best of our knowledge, this paper is the first to identify that this assumption does not hold in attention mechanisms due to the differing projection weights for query and key vectors. This OOD characteristic of query vectors compromises the efficiency gains anticipated from ANNS, resulting in an increased demand for memory access. Our empirical analysis reveals that to maintain acceptable accuracy levels, it is imperative to scan 30% of all key vectors.
@@
In this work, we present RetrievalAttention, an efficient method for accelerating long-context LLM generation. RetrievalAttention employs dynamic sparse attention during token generation, allowing the most critical tokens to emerge from the extensive context data. To address the OOD issue, RetrievalAttention constructs a vector index tailored for the attention mechanism, focusing on the distribution of queries rather than key similarities. This approach allows for traversal of only a small subset of key vectors (1% to 3%), effectively identifying the most relevant tokens to achieve accurate attention scores and results. To optimize resource utilization, RetrievalAttention retains KV vectors in the GPU memory following static patterns (e.g., StreamingLLM [10]), while offloading the majority of KV vectors to CPU memory for index construction. During token generation, RetrievalAttention efficiently retrieves critical tokens using vector indexes on the CPU and merges the partial attention results from both the CPU and GPU. This strategy enables RetrievalAttention to perform attention computation with reduced latency and minimal GPU memory utilization.
@@
We evaluate RetrievalAttention for accuracy and efficiency on both commodity GPUs (4090) and high-end GPUs (A100) on three long-context LLMs across various long-context benchmarks like $\infty$-Bench [12] and RULER [13]. For the 128K context on the 4090 GPU, RetrievalAttention achieves $4.9\times$ and $1.98\times$ decoding-latency reduction compared to the retrieval method based on exact KNN and traditional ANNS indexing respectively, while maintaining the same accuracy as full attention. To the best of our knowledge, RetrievalAttention is the first solution that supports running 8B-level models on a single 4090 GPU (24GB) with acceptable latency and almost no accuracy degradation.
@@ 2.3 Dynamic and Sparse Attention
Table 1: Decoding latency (on single A100) and GPU memory required for KV Cache of Llama-2-7B.
  Prompt Length 16k 64k 128k 256k 512k 1M
  Total Latency (s) 1.5 10.2 32.8 111 465 1,765
|-           FFN (s) 0.6 3.1 7.6 15 31 70
|-   Attention (s) 0.9 7.1 25.2 96 434 1,695
GPU Memory
KV Cache (GB)
7.8 31.3 62.5 125 250 500
@@
Despite the large size of the context, only a small proportion of tokens actually dominate the generation accuracy. Figure 2(a) shows the distribution of $|a_{t,i}|$ in Equation 1 for a query vector from Llama-2-7B with a prompt length of 64,000 tokens. We observe that the top 500 tokens dominate the values of $|a_{t,i}|$, while the remaining tokens are approximately zero. A high attention score indicates that two vectors are close, as measured by the inner product. Therefore, the sparsity of attention scores means that the relevant keys to the query are very few. We measure the mean-square-error (MSE) of the attention output if we use the top-$k$ $|a_{t,i}|$ as an approximation. We find that it only needs 36 tokens to achieve a very low MSE ($10^{-6}$) of the full attention, showing a high sparsity ratio ($99.9\%$).
Moreover, we observe that as the LLM continues generating new tokens, critical key vectors are dynamically changing. This means that tokens considered important in previous queries may not be critical in subsequent queries, and vice versa. The dynamic sparsity shows a promising path to approximately compute attention with greatly reduced cost and without sacrificing the model accuracy. For each query, if we can accurately determine the relevant key-value vectors with higher importance, minimum GPU memory and a much lower time complexity can be achieved for attention computation.
@@ 3.1 Approximated Attention
Building upon the full attention defined in Equation 1, RetrievalAttention approximates the full attention output $\mathbf{o}_{t}$ by selectively utilizing the KV vectors associated with high attention scores (i.e., $a_{t,i}$). Specifically, we define $\mathcal{I}_{t,\epsilon}$ as a subset of token indices for which the attention score surpasses $\epsilon$. Consequently, a sparse attention mechanism, which only considers tokens located in $\mathcal{I}_{t,\epsilon}$, can be defined as follows:
@@ 4.1 Experimental Setup
Testbed, Models, and Configurations. We conduct experiments on a server equipped with one NVIDIA RTX 4090 GPU (24GB memory) and an Intel i9-10900X CPU with 20 cores and 128GB DRAM. The experiment results using NVIDIA A100 GPU can be found in §A.4. We implement RetrievalAttention on three state-of-the-art long-context LLMs, including Llama-3-8B-Instruct-262k[24], Yi-6B-200K[25], and Yi-9B-200K[26]. We set batch_size=1batch\_size=1 to ensure the size of indexes and KV vectors do not exceed the capacity of the CPU memory in long contexts.
Baselines. We compare RetrievalAttention with the following training-free baselines. (1) Full attention without KV cache as well as the version with KV cache using vLLM [27]. (2) StreamingLLM[10]: it retains initial tokens along with fixed-length recent tokens in the GPU memory and discards remaining tokens. (3) SnapKV[11]: it caches the active tokens observed from the last window of the prompts. (4) InfLLM[6]: it separates KV cache of continuous token sequences into blocks and select representative vectors for each block. In the decoding phase, the current query scans all representative vectors and retrieve top-$k$ blocks with the highest similarity. We do not include Quest [5] for evaluation because its implementation ***https://github.com/mit-han-lab/quest does not support the GQA model.
@@ 5 Related Works
FlexGen [4] and Lamina [33] offload the KV cache to CPU memory, but they struggle with slow and costly full-attention computation. By identifying the dynamic nature of important KV vectors for different queries, recent work chooses to retain all of the KV cache and dynamically attend to different parts of KV vectors based on the current query. Quest [5] partitions the KV cache into blocks and selects a representative key vector for each block. For a given query, it scans all representative key vectors and attends top-$k$ blocks with the highest attention scores. InfLLM [6] adopts a similar strategy as Quest but offloads most KV cache blocks to the CPU memory to support longer contexts. Due to block-based organization and retrieval, the accuracy of representative vectors significantly impacts the effectiveness of those methods for obtaining important tokens. SparQ [7], InfiniGen [8], and LoKi [9] approximate the most relevant top-$k$ keys corresponding to a given query by reducing the head dimension. RetrievalAttention instead organizes the KV cache using ANNS indexes, allowing the retrieval of important tokens with high recalls and low cost. The concurrent work MagicPiG [34] employs locality-sensitive-hashing (LSH) to retrieve critical KV vectors. However, it fails to address the out-of-distribution issue between query and key vectors, necessitating retrieving a large portion of KV cache to maintain the model accuracy.
@@ 6 Conclusion
We propose RetrievalAttention, a method that offloads most KV vectors to CPU memory and leverages vector search for dynamic sparse attention to minimize inference cost. RetrievalAttention identifies the different distributions of the query and key vectors and employs a attention-aware approach to efficiently find critical tokens for model generation. Experimental results demonstrate RetrievalAttention effectively achieves $4.9\times$ and $1.98\times$ decoding speedup than exact KNN and traditional ANNS methods, on a single RTX4090 GPU for a context of 128k tokens. RetrievalAttention is the first system that supports running 8B-level LLMs with 128K tokens on a single 4090 (24GB) GPU with an acceptable latency cost and without compromising model accuracy.
@@ A.4 Decoding Latency on A100
We test the generality of RetrievalAttention by measuring its performance on a server with one A100 (80GB) and one AMD EPYC CPU with 24 cores. We show the token-generation latency of different methods on three models in Table 8. Since the KV cache of full attention is disabled, all prompt tokens need to be recalculated during the decoding, incurring a very high decoding latency. By enabling the KV cache with the PageAttention optimization in vLLM, the decoding latency is significantly reduced. However, vLLM suffers from OOM issue with the increase of context length, which we elaborate further later. Other KV cache dropping or block retrieval methods including StreamingLLM, SnapKV, and InfLLM achieve faster decoding speed, but this is at the expense of a significant drop in model accuracy. In contrast, RetrievalAttention does not compromise generation accuracy while achieving much lower decoding latency than IVF and Flat because of the efficient mitigation of out-of-distribution problem.
@@ Appendix D Additional Related Work
Since the quadratic complexity of attention has become the bottleneck of LLM efficiency for long context applications, numerous works have studied to design sparse transformers to reduce the computational and memory complexity of the self-attention mechanism. Some works restrict the attention computation to predefined patterns including sliding windows [36], dilated windows [37], or mixture of different patterns [38, 39]. Some approaches use cluster-based sparsity based on hash value [40] or KNN algorithms [41, 42]. All these solutions require pre-training a model from scratch, which do not work for our target to out-of-box usage of LLMs. Although some approaches [6, 7] exploit the dynamic sparse nature of LLMs, they often use some estimation using low-rank hidden states or post-statistical approaches, which incurs high overhead but with low accuracy. Moreover, all these approaches have to maintain full KV vectors on GPU with only accelerated inference by reduced memory movement, which does not solve the challenge of commodity GPUs with limited GPU memory.

### v2

@@ Abstract
Transformer-based Large Language Models (LLMs) have become increasingly important. However, due to the quadratic time complexity of attention computation, scaling LLMs to longer contexts incurs extremely slow inference latency and high GPU memory consumption for caching key-value (KV) vectors. This paper proposes RetrievalAttention, a training-free approach to both accelerate attention computation and reduce GPU memory consumption. By leveraging the dynamic sparsity of attention mechanism, RetrievalAttention proposes to use approximate nearest neighbor search (ANNS) indexes for KV vectors in CPU memory and retrieves the most relevant ones with vector search during generation. Unfortunately, we observe that the off-the-shelf ANNS indexes are often ineffective for such retrieval tasks due to the out-of-distribution (OOD) between query vectors and key vectors in attention mechanism. RetrievalAttention addresses the OOD challenge by designing an attention-aware vector search algorithm that can adapt to the distribution of query vectors. Our evaluation shows that RetrievalAttention only needs to access 1–3% of data while maintaining high model accuracy. This leads to significant reduction in the inference cost of long-context LLMs with much lower GPU memory footprint. In particular, RetrievalAttention only needs a single NVIDIA RTX4090 (24GB) for serving 128K tokens in LLMs with 8B parameters, which is capable of generating one token in 0.188 seconds.
@@ 1 Introduction
Recent transformer-based Large Language Models [1] have shown remarkable capabilities in processing long contexts. For instance, Gemini 1.5 Pro [2] has supported the context window of up to 10 million tokens. While this is promising for analyzing extensive data, supporting longer context windows also introduces challenges for inference efficiency due to the quadratic complexity of attention computation. To enhance efficiency, KV caching, a technique that retains past Key and Value vectors, has been widely adopted to prevent redundant computations. However, KV caching-based systems face two primary issues: (a) substantial GPU memory requirements particularly for long contexts, e.g., the Llama-2-7B model requires approximately 500GB per million tokens in FP16 format; and (b) inference latency increases linearly to the context size, primarily due to the time needed to access cached tokens — a common issue across various computing devices, including GPUs. Therefore, reducing token access and storage costs is vital for enhancing inference efficiency.
@@
The solution lies in leveraging the dynamic sparsity inherent in the attention mechanism [3]. This refers to the phenomenon where each query vector significantly interacts with only a limited subset of key and value vectors, with the selection of these tokens varying dynamically based on individual queries. Prior work [4, 5, 6, 7, 8, 9] has proposed various techniques to capitalize on this observation to improve the efficiency of attention computation. However, most of these methods identify important tokens either statically [10, 11] or heuristically [6, 7, 5], leading to imprecise approximations that often result in a significant performance drop. We observe that the Approximate Nearest Neighbor Search (ANNS) index is particularly effective in this context. When using the inner product as the similarity measurement, ANNS precisely aligns with the attention mechanism, represented as $\mathbf{q}\mathbf{K}^{T}$. It can directly identify the most critical key vectors with maximum inner product value, yielding a higher accuracy compared to previous static or heuristic methods (as illustrated in Figure 1).
Figure 1: RetrievalAttention achieves similar task accuracy as full attention but exhibits extremely low decoding latency.
Prompt Length 128K 256K 512K 1M
Total Latency (s) 32.8 111 465 1,765
             FFN (s) 7.6 15 31 70
     Attention (s) 25.2 96 434 1,695
GPU Memory
KV Cache (GB)
62.5 125 250 500
Table 1: Decoding latency (on one A100) and GPU memory required for KV cache of Llama-2-7B across different context lengths.
@@
Leveraging ANNS for attention mechanisms presents a unique challenge: the out-of-distribution (OOD) problem between query and key vectors. Most ANNS engines operate under the assumption that both query and key vectors are drawn from the same distribution. However, due to the different projection weights for query and key vectors in attention mechanism, this assumption often does not hold in this context. Unfortunately, the effectiveness of ANNS degrades significantly if the query vectors are OOD. In particular, our empirical analysis shows that maintaining an acceptable accuracy level for inference quality can require scanning 30% of all key vectors to identify the critical ones, which fails to exploit the inherent sparsity in attention mechanism. To the best of our knowledge, this is the first work that identifies the challenge of OOD in using ANNS for attention computation.
@@
In this work, we present RetrievalAttention, an efficient method for accelerating long-context LLM generation. RetrievalAttention employs dynamic sparse attention during token generation, allowing the most critical tokens to emerge from the extensive context data. To address the challenge of OOD, RetrievalAttention proposes a vector index tailored for the attention mechanism, focusing on the distribution of queries rather than key similarities. This approach allows for traversal of only a small subset of key vectors (1% to 3%), effectively identifying the most relevant tokens to achieve accurate attention scores and inference results. In addition, RetrievalAttention reduces GPU memory consumption by retaining a small number of KV vectors in the GPU memory following static patterns (e.g., similar to StreamingLLM [10]) and offloading the majority of KV vectors to CPU memory for index construction. During token generation, RetrievalAttention efficiently retrieves critical tokens using vector indexes on the CPU and merges the partial attention results from both the CPU and GPU. This strategy enables RetrievalAttention to perform attention computation with reduced latency and minimal GPU memory footprint.
@@
We evaluate the accuracy and efficiency of RetrievalAttention on both commodity GPUs (4090) and high-end GPUs (A100) on three long-context LLMs across various long-context benchmarks like $\infty$-Bench [12] and RULER [13]. For the 128K context on the 4090 GPU, RetrievalAttention achieves $4.9\times$ and $1.98\times$ decoding-latency reduction compared to the retrieval method based on exact KNN and traditional ANNS indexing respectively, while maintaining the same accuracy as full attention. To the best of our knowledge, RetrievalAttention is the first solution that supports running 8B-level models on a single 4090 GPU (24GB) with acceptable latency and almost no accuracy degradation.
@@ 2.3 Dynamic and Sparse Attention
Despite the large size of the context, only a small proportion of tokens actually dominate the generation accuracy. Figure 2(a) shows the distribution of $|a_{t,i}|$ in Equation 1 for a query vector from Llama-2-7B with a prompt of 64,000 tokens. We observe that the top 500 tokens dominate the values of $|a_{t,i}|$, while the remaining tokens contribute approximately zero. A high attention score indicates that two vectors are close, as measured by the inner product. Therefore, the sparsity of attention scores means that the relevant keys to the query are very few. We measure the mean-square-error (MSE) of the attention output if we use the top-$k$ $|a_{t,i}|$ as an approximation. We find that it only needs 36 tokens to achieve a very low MSE ($10^{-6}$) of the full attention, showing a high sparsity ratio ($99.9\%$).
Moreover, we observe that as the LLM continues generating new tokens, critical key vectors change dynamically. This means that tokens considered important in previous queries may not be critical in subsequent queries, and vice versa. The dynamic sparsity shows a promising path to approximately compute attention with greatly reduced cost and without sacrificing the model accuracy. For each query, if we can accurately determine the relevant key-value vectors with higher importance, minimum GPU memory and a much lower time complexity can be achieved for attention computation.
@@ 3.1 Approximated Attention
Based on the Equation 1, RetrievalAttention approximates the full attention output $\mathbf{o}_{t}$ by selectively utilizing the KV vectors associated with high attention scores (i.e., $a_{t,i}$). Specifically, we define $\mathcal{I}_{t,\epsilon}$ as a subset of token indices for which the attention score surpasses $\epsilon$. Consequently, a sparse attention mechanism, which only considers tokens located in $\mathcal{I}_{t,\epsilon}$, can be defined as follows:
@@ 4.1 Experimental Setup
Testbed, Models, and Configurations. We conduct experiments on a server equipped with one NVIDIA RTX 4090 GPU (24GB memory) and an Intel i9-10900X CPU with 20 cores and 128GB DRAM. The experiment results using NVIDIA A100 GPU can be found in §A.4. We implement RetrievalAttention on three state-of-the-art long-context LLMs, including Llama-3-8B-Instruct-262k[24], Yi-6B-200K[25], and Yi-9B-200K[26]. To show practical speedup of RetrievalAttention and ensure the CPU memory consumption in long contexts do not exceed the capacity of DRAM, we follow previous work [5] to run the benchmark in real-world single-batch scenarios.
Baselines. We compare RetrievalAttention with the following training-free baselines. (1) Full attention without KV cache as well as the version with KV cache using vLLM [27]. (2) StreamingLLM[10]: it retains initial tokens along with fixed-length recent tokens in the GPU memory and discards remaining tokens. (3) SnapKV[11]: it caches the active tokens observed from the last window of the prompts. (4) InfLLM[6]: it separates KV cache of continuous token sequences into blocks and select representative vectors for each block. In the decoding phase, the current query scans all representative vectors and retrieve top-$k$ blocks with the highest similarity. We do not include Quest [5] for evaluation because its implementation  [^1] https://github.com/mit-han-lab/quest does not support the GQA model.
@@ 5 Related Works
FlexGen [4] and Lamina [33] offload the KV cache to CPU memory, but they struggle with slow and costly full-attention computation. By identifying the dynamic nature of important KV vectors for different queries, recent work chooses to retain all of the KV cache and dynamically attend to different parts of KV vectors based on the current query. Quest [5] partitions the KV cache into blocks and selects a representative key vector for each block. For a given query, it scans all representative key vectors and attends top-$k$ blocks with the highest attention scores. InfLLM [6] adopts a similar strategy as Quest but offloads most KV cache blocks to the CPU memory to support longer contexts. Due to block-based organization and retrieval, the accuracy of representative vectors significantly impacts the effectiveness of those methods for obtaining important tokens. SparQ [7], InfiniGen [8], and LoKi [9] approximate the most relevant top-$k$ keys corresponding to a given query by reducing the head dimension. RetrievalAttention instead organizes the KV cache using ANNS indexes, allowing the retrieval of important tokens with high recalls and low cost. The concurrent work MagicPiG [34] and PQCache [35] employ locality-sensitive-hashing (LSH) and PQ centroids respectively to retrieve critical tokens. However, they fail to address the OOD issue in attention, necessitating retrieving a large portion of KV cache (e.g., 20%) to maintain the model accuracy.
@@ 6 Conclusion
We propose RetrievalAttention, a method that offloads most KV vectors to CPU memory and leverages vector search for dynamic sparse attention to minimize inference cost. RetrievalAttention identifies the different distributions of the query and key vectors and employs an attention-aware approach to efficiently find critical tokens for model generation. Experimental results demonstrate that RetrievalAttention effectively achieves $4.9\times$ and $1.98\times$ decoding speedup than exact KNN and traditional ANNS methods, on a single RTX4090 GPU for a context of 128K tokens. RetrievalAttention is the first system that supports running 8B-level LLMs with 128K tokens on a single 4090 (24GB) GPU with an acceptable latency cost and without compromising model accuracy.
@@ A.4 Decoding Latency on A100
We test the generality of RetrievalAttention by measuring its performance on a server with one A100 (80GB) and one AMD EPYC CPU with 24 cores and 220GB DRAM. We show the token-generation latency of different methods on three models in Table 8. Since the KV cache of full attention is disabled, all prompt tokens need to be recalculated during the decoding, incurring a very high decoding latency. By enabling the KV cache with the PageAttention optimization in vLLM, the decoding latency is significantly reduced. However, vLLM suffers from OOM issue with the increase of context length, which we elaborate further later. Other KV cache dropping or block retrieval methods including StreamingLLM, SnapKV, and InfLLM achieve faster decoding speed, but this is at the expense of a significant drop in model accuracy. In contrast, RetrievalAttention does not compromise generation accuracy while achieving much lower decoding latency than IVF and Flat because of the efficient mitigation of out-of-distribution problem.
@@ Appendix D Additional Related Work
Since the quadratic complexity of attention has become the bottleneck of LLM efficiency for long context applications, numerous works have studied to design sparse transformers to reduce the computational and memory complexity of the self-attention mechanism. Some works restrict the attention computation to predefined patterns including sliding windows [37], dilated windows [38], or mixture of different patterns [39, 40]. Some approaches use cluster-based sparsity based on hash value [41] or KNN algorithms [42, 43].These solutions either require pre-training a model from scratch or target limited scenarios like CPU-only, which do not work for our target to out-of-box usage of LLMs on the GPU-CPU architecture. Although some approaches [6, 7] exploit the dynamic sparse nature of LLMs, they often use some estimation using low-rank hidden states or post-statistical approaches, which incurs high overhead but with low accuracy. Moreover, all these approaches have to maintain full KV vectors on GPU with only accelerated inference by reduced memory movement, which does not solve the challenge of commodity GPUs with limited GPU memory.